How many time constants (a decimal number) must elapse before a capacitor in a series RC circuit is charged to 88.0% of its equilibrium charge?
Answer #1

output = 1 - e^-(t/T) = .88

ln .18 = - t/T

t = -T ln .18 = 1.7 T